# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 3.3 States
for $i=1,2$. Using Prop. 3.7 and Lemma 3.6 and the fact that the commutator of two derivations that preserve $\psi$ preserves $\psi$, one checks that $\tilde{\gamma}$ takes ${\mathfrak{D}}_{al}^{\psi}(U\cup V)$ to ${\mathfrak{D}}_{al}^{\psi}(U)\oplus{\mathfrak{D}}_{al}^{\psi}(V)$. Using Prop. 3.7 $iii)$ and the fact that $\gamma_{1}^{U,V}({\mathsf{F}})+\gamma_{2}^{U,V}({\mathsf{F}})={\mathsf{F}}$ and $\gamma_{1}^{U^{\prime},V^{\prime}}({\mathsf{H}})+\gamma_{2}^{U^{\prime},V^{\prime}}({\mathsf{H}})={\mathsf{H}}$, we get $\tilde{\gamma}_{1}({\mathsf{F}})+\tilde{\gamma}_{2}({\mathsf{F}})={\mathsf{F}}$, as desired. ∎
We showed that one can attach Fréchet-Lie algebras ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}^{\psi}(U)$ to any $U\subset{\mathbb{R}}^{n}$. These form a pre-cosheaf over the pre-ordered set of subsets of ${\mathbb{R}}^{n}$. Moreover, Proposition 3.5, Corollary 3.1, and (when $\psi$ is gapped) Proposition 3.8 show that these spaces satisfy the cosheaf condition and Property I for transverse pairs $U,V\subset{\mathbb{R}}^{n}$. What prevents the functors ${\mathfrak{D}}_{al}$ and ${\mathfrak{D}}^{\psi}_{al}$ from forming local Lie algebras is the fact that pairs of subsets $U,V\subset{\mathbb{R}}^{n}$ generally do not intersect transversely. This problem can be resolved by restricting to a suitable set of subsets of ${\mathbb{R}}^{n}$ that have well-behaved intersections. In the next section we identify one such set, the set of semilinear subsets, prove that they form a Grothendieck site, and discuss some properties of this site.
@@ 4.1 Semilinear sets and their thickenings
A semilinear set in ${\mathbb{R}}^{n}$ is a subset of ${\mathbb{R}}^{n}$ which can be defined by means of a finite number of linear equalities and strict linear inequalities. More precisely, a basic semilinear set in ${\mathbb{R}}^{n}$ is an intersection of a finite number of hyperplanes and open half-spaces, and a semilinear set is a finite union of basic semilinear sets. The set of semilinear subsets of ${\mathbb{R}}^{n}$ will be denoted ${\mathcal{S}}_{n}$. Projections ${\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ map ${\mathcal{S}}_{m+n}$ to ${\mathcal{S}}_{m}$ [23]. A map ${\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ is called semilinear iff its graph is a semilinear subset of ${\mathbb{R}}^{m+n}$. The composition of two semilinear maps is a semilinear map [23].
@@
Recall that for any set $U\subset{\mathbb{R}}^{n}$, we write $U^{r}:=\{x\in{\mathbb{R}}^{n}:\exists y\in U\text{ s.t }d(x,y)\leq r\}$ and call this the $r$-thickening of $U$. It is easy to see that if $U$ is closed, then $U^{r}$ is also closed for any $r\geq 0$.
Lemma 4.2.
@@
$\Delta_{r}=\{(x,y)\in{\mathbb{R}}^{2n}|d(x,y)\leq r\}.$
By the previous lemma, $\Delta_{r}$ is semilinear. On the other hand, $U^{r}$ is the projection to the first ${\mathbb{R}}^{n}$ of $\Delta_{r}\cap({\mathbb{R}}^{n}\times U)\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}$. Since intersection and projection preserve the set of semilinear sets, the lemma is proved. ∎
Lemma 4.3.
@@
Suppose $x,y\in U^{r}$, and suppose $x^{\prime},y^{\prime}\in U$ satisfy $d(x,x^{\prime})\leq r$ and $d(y,y^{\prime})\leq r$. Then for any $t\in[0,1]$ we have $d(tx+(1-t)y,tx^{\prime}+(1-t)x^{\prime})\leq t^{2}d(x,x^{\prime})+(1-t)^{2}d(y,y^{\prime})\leq r.$
A polyhedron in ${\mathbb{R}}^{n}$ is an intersection of a finite number of closed half-spaces. A polyhedron is closed, but not necessarily compact. A closed semilinear set is the same as a finite union of polyhedra. Conversely, according to Theorem 19.6 from [24], a polyhedron can be described as a closed convex semilinear set. Combining this with the above lemmas, we get
Corollary 4.1.

### v3

@@ 2.3 States
for $i=1,2$. Using Prop. 49 and Lemma 52 and the fact that the commutator of two derivations that preserve $\psi$ preserves $\psi$, one checks that $\tilde{\gamma}$ takes ${\mathfrak{D}}^{\psi}_{al}(U\cup V)$ to ${\mathfrak{D}}^{\psi}_{al}(U)\oplus{\mathfrak{D}}^{\psi}_{al}(V)$. Using Prop. 49 $iii)$ and the fact that $\gamma_{1}^{U,V}({\mathsf{F}})+\gamma_{2}^{U,V}({\mathsf{F}})={\mathsf{F}}$ and $\gamma_{1}^{U^{\prime},V^{\prime}}({\mathsf{H}})+\gamma_{2}^{U^{\prime},V^{\prime}}({\mathsf{H}})={\mathsf{H}}$, we get $\tilde{\gamma}_{1}({\mathsf{F}})+\tilde{\gamma}_{2}({\mathsf{F}})={\mathsf{F}}$, as desired. ∎
Example 53.
Consider a Fock state $\omega$ of a CAR algebra, Example 48, corresponding to a projector $P$ on $\ell^{2}(\Lambda,H)$. Let $U_{i}$, $i\in I$, be a finite set of mutually transverse closed sets such that $\cup_{i}U_{i}={\mathbb{R}}^{n}$. This gives a covering of ${\mathbb{R}}^{n}$. Let $\mathsf{Q}$} be the generator of the canonical $U(1)$ action on the CAR algebra, Example 6. Since $\omega$ is $U(1)$-invariant, ${\mathsf{Q}}\in{\mathfrak{D}}_{al}^{\omega}({\mathbb{R}}^{n})$, hence by Prop. 50 there exist ${\mathsf{Q}}_{i}\in{\mathfrak{D}}_{al}^{\omega}(U_{i})$ such that $\sum_{i}{\mathsf{Q}}_{i}={\mathsf{Q}}$. In this special case one can exhibit such ${\mathsf{Q}}_{i}$ explicitly. For every $i\in I$, let $\chi_{i}:\Lambda\rightarrow\{0,1\}$ be the indicator function of $U_{i}$. For every $x\in\Lambda$ let $n_{x}=\sum_{i\in I}\chi_{i}(x)$. We define the functions $f_{i}:\Lambda\rightarrow{\mathbb{R}}$, $i\in I$, by
$f_{i}(x)=\frac{\chi_{i}(x)}{n_{x}},\quad\forall x\in\Lambda,$ (97)
so that $\sum_{i}f_{i}=1$. We can regard each $f_{i}$ as a bounded operator on $\ell^{2}(\Lambda,H)$. Now we let ${\mathsf{Q}}_{i}$ be the derivation ${\mathsf{F}}(h_{i})$ (see Example 19) corresponding to the following operators $h_{i}$ on $\ell^{2}(\Lambda,H)$:
$h_{i}=Pf_{i}P+(1-P)f_{i}(1-P).$ (98)
It is easy to see that $\sum_{i\in I}h_{i}=1,$ hence $\sum_{i\in I}{\mathsf{F}}(h_{i})={\mathsf{F}}(1)={\mathsf{Q}}$. One can also check that ${\mathsf{Q}}_{i}\in{\mathfrak{D}}_{al}^{\omega}(U_{i})$.
3 Local Lie systems
In this section we define the central notion of this paper: local Lie systems over a site. Physically, this notion formalizes what physicists call gauge Lie algebras. Mathematically, a local Lie system is an algebraic object that encodes the compatibility conditions between the spaces ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}^{\psi}_{al}(U)$ attached to regions $U\subseteq\mathbb{R}^{n}$. That is, local Lie systems are an axiomatization of Lemma 24, Proposition 29 (resp. Proposition 50), and Corollary 34 for the spaces ${\mathfrak{D}}_{al}(U)$ (resp. ${\mathfrak{D}}^{\psi}_{al}(U)$). Our axiomatization involves a pre-cosheaf of Lie algebras over a suitable category of subsets of the space. The space can be either a discrete metric space, as in this paper, or a manifold, as in field theory. A basic construction which applies to any pre-cosheaf of vector spaces is the Čech functor. We show in subsection 3.2 that in the case of local Lie systems it produces a Differential Graded Lie algebra (DGLA). This DGLA will be used later on to construct topological invariants of gapped states.
3.1 Locality and (pre-)cosheaves
As an example, consider the Lie algebra of gauge transformations, i.e. the Lie algebra ${\mathfrak{G}}(M):=C^{\infty}(M,{\mathfrak{g}})$ of smooth functions on $M$ with values in a finite-dimensional Lie algebra $\mathfrak{g}$}. It is a global object attached to $M$. To “localize” it, for any open $U\subseteq M$ we define the Lie algebra ${\mathfrak{G}}(U)$ to be the space of smooth $\mathfrak{g}$}-valued functions on $M$ whose closed support is contained in $U$. In particular, ${\mathfrak{G}}(\varnothing)=0$. For any inclusion of opens $U\subseteq V$ we have a homomorphism of Lie algebras $\iota_{VU}:{\mathfrak{G}}(U)\rightarrow{\mathfrak{G}}(V)$ such that the Lie algebras ${\mathfrak{G}}(U)$ assemble into a pre-cosheaf $\mathfrak{G}$} of Lie algebras on $M$. This is a coflasque pre-cosheaf, i.e. all its structure maps $\iota_{VU}$ are injective. [^6] The terminology comes from sheaf theory, where a pre-sheaf ${\mathcal{F}}:Open(M)^{opp}\rightarrow{\mathcal{C}}$ is called flasque if for any $U\subseteq V$ the restriction morphism ${\mathcal{F}}(V)\rightarrow{\mathcal{F}}(U)$ is a surjection.
Here the first arrow is $\iota_{U,U\cap V}\oplus(-\iota_{V,U\cap V})$ and the second arrow is $\iota_{U\cup V,U}\oplus\iota_{U\cup V,V}$. Exactness follows from the existence of a partition of unity for the cover ${\mathfrak{U}}=\{U,V\}$ of $U\cup V$. In words, the exactness of the sequence (99) means that any element of ${\mathfrak{G}}(U\cup V)$ can be decomposed as a sum of elements of subalgebras attached to $U$ and $V$ modulo ambiguities which take values in the subalgebra attached to $U\cap V$.
More generally, the existence of a partition of unity implies that for any finite collection of opens $U_{i}$, $i\in I$ the following sequence is exact [12]:
For a compact $M$, this implies that the pre-cosheaf of vector spaces $\mathfrak{G}$} is a cosheaf of vector spaces. The cosheaf property is a compatibility of the pre-cosheaf with the notions of intersection and union of opens and expresses a stronger form of locality.
Note that the maps in the above exact sequences are not Lie algebra homomorphisms. Hence $\mathfrak{G}$} is not a cosheaf of Lie algebras. Nevertheless, the following additional property of $\mathfrak{G}$} can be regarded as a form of locality of the Lie bracket.
Definition 54.
A coflasque pre-cosheaf of Lie algebras $\mathfrak{F}$} is said to satisfy Property I if for any two opens $U,V$ one has ${\mathfrak{F}}(U),{\mathfrak{F}}(V)]\subseteq{\mathfrak{F}}(U\cap V)$.
Clearly, $\mathfrak{G}$} satisfies Property I. In particular, elements of ${\mathfrak{G}}(M)$ that are supported on non-intersecting opens commute.
In the previous section, for every quantum lattice system with a gapped state $\psi$ we defined Lie algebras ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}^{\psi}_{al}(U)$ attached to subsets in ${\mathbb{R}}^{n}$ in such a way that ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(U^{r})$ and ${\mathfrak{D}}^{\psi}_{al}(U)={\mathfrak{D}}^{\psi}_{al}(U^{r})$ for any $U$ and any $r\geq 0$. Intuitively, this means that localization on the lattice is only approximate. To phrase locality of symmetries in lattice systems, one needs to replace the set $Open(M)$ of open subsets of $M$ with a more general structure which admits the notions of intersection, union, and cover.
For any pre-ordered set $(X,\leq)$ there is a natural notion of intersection and union. The intersection of $U,V\in X$ can be defined as the greatest lower bound (or meet) of both $U$ and $V$, i.e. a $W\in X$ such that $W\leq U$, $W\leq V$, and for any $W^{\prime}\leq U,V$ we have $W^{\prime}\leq W$. The meet of $U$ and $V$ is denoted $U\wedge V$. Similarly, the union of $U$ and $V$ can be defined as the smallest upper bound (or join) of both $U$ and $V$. It is denoted $U\vee V$. For a general pre-ordered set, the meet and join may not exist for all pairs of objects. If they exist, they are unique up to isomorphism. We will assume that $(X,\leq)$ is such that $U\wedge V$ and $U\vee V$ exist for all $U,V\in X$. [^7] If we turn $X$ into a poset by identifying isomorphic objects, then this means that the poset is a lattice in the sense of order theory. The existence of all pairwise meets and joins implies the existence of all finite meets and joins. In the case of the pre-ordered set $Open(M)$ arbitrary (i.e. not necessarily finite) joins make sense.
Finally, to define covers of elements of $X$, let us assume that $U\wedge(V\vee W)\leq(U\wedge V)\vee(U\wedge W)$ for all $U,V,W\in X$. [^8] The opposite relation is automatic, so this condition ensures that $U\wedge(V\vee W)\simeq(U\wedge V)\vee(U\wedge W)$ for all $U,V,W\in X$. This is equivalent to saying that the poset corresponding to $X$ is a distributive lattice. We will say that $X$ is a distributive pre-ordered set. This condition is certainly satisfied for $Open(M)$. We say that a collection ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ of elements of $X$ covers $A\in X$ iff $U_{i}\leq A$ for all $i\in I$ and $A\leq\bigvee_{i\in I}U_{i}$. This definition ensures that if $\mathfrak{U}$} covers $A$, then for any $B\leq A$ the collection ${\mathfrak{U}}\wedge B=\{U_{i}\wedge B\}_{i\in I}$ covers $B$.
Definition 55.
A local Lie system over $(X,\leq)$ is a coflasque pre-cosheaf of Lie algebras with Property I which is also a cosheaf of vector spaces over the corresponding site. A morphism of local Lie systems is a morphism of the underlying pre-cosheaves of Lie algebras.
Remark 56.
Let $\mathfrak{F}$} be a local Lie system. Then for any $U\leq V$ the Lie algebra ${\mathfrak{F}}(U)$ is an ideal in ${\mathfrak{F}}(V)$. One can equivalently define a local Lie system as a pre-cosheaf of Lie algebras over $(X,\leq)$ which is a cosheaf of vector spaces and such that all co-restriction maps are inclusions of Lie ideals.
Remark 57.
In this paper all Lie algebras will be Fréchet-Lie algebras and all morphisms will be continuous. We define a local Fréchet-Lie system over $(X,\leq)$ to be a coflasque pre-cosheaf of Fréchet-Lie algebras with Property I which is also a cosheaf of vector spaces.
Lemma 58.
Let $X$ be a distributive pre-ordered set with finite meets $\wedge$ and joins $\vee$. Write $A\cong B$ to mean $A\leq B$ and $B\leq A$. Equip $X$ with the Grothendieck topology of finite covers: a finite family $\{U_{i}\}_{i\in I}$ is a cover of $U$ if $U_{i}\leq U$ for all $i\in I$ and $U\leq\bigvee_{i\in I}U_{i}$. Let $\mathfrak{F}$} be a pre-cosheaf of vector spaces on $X$, with structure maps $\iota_{VU}:{\mathfrak{F}}(U)\to{\mathfrak{F}}(V)$ for $U\leq V$.
Then $\mathfrak{F}$} is a cosheaf for the above topology if and only if for every $U,V\in X$ the sequence
${\mathfrak{F}}(U\wedge V)\xrightarrow{\ \alpha\ }{\mathfrak{F}}(U)\oplus{\mathfrak{F}}(V)\xrightarrow{\ \beta\ }{\mathfrak{F}}(U\vee V)\to 0$ (101)
is exact, where
$\alpha(t)=\bigl(\iota_{U,\,U\wedge V}(t),\,-\iota_{V,\,U\wedge V}(t)\bigr),\qquad\beta(a,b)=\iota_{U\vee V,\,U}(a)+\iota_{U\vee V,\,V}(b).$ (102)
The following proof is essentially a restatement of the proof of Proposition 1.3 in [12], adapted to the site $X$:
Proof.
The “only if” direction is the cosheaf axiom for the $2$-cover $\{U,V\}$ of $U\vee V$.
Conversely, assume exactness for all pairs. Since $\mathfrak{F}$} sends $\cong$ to isomorphisms, for a cover $\{U_{i}\leq W\}$ one may replace $W$ by $\bigvee_{i}U_{i}$. Thus it suffices to treat a finite family $U_{1},\dots,U_{n}$ with $U:=\bigvee_{i=1}^{n}U_{i}$.
Consider the standard sequence
$\bigoplus_{i<j}{\mathfrak{F}}(U_{i}\wedge U_{j})\xrightarrow{\ \alpha\ }\bigoplus_{i=1}^{n}{\mathfrak{F}}(U_{i})\xrightarrow{\ \beta\ }{\mathfrak{F}}(U)\to 0,$ (103)
with $\beta((s_{i}))=\sum_{i}\iota_{U,U_{i}}(s_{i})$ and $\alpha$ the alternating sum on pairwise overlaps. Surjectivity of $\beta$ follows by induction on $n$ using the exactness for the pair $\bigl(\bigvee_{i=1}^{n-1}U_{i},\;U_{n}\bigr)$ and the inductive hypothesis for $\bigvee_{i=1}^{n-1}U_{i}$.
For exactness in the middle, let $(s_{i})\in\ker\beta$ and set $V:=\bigvee_{i=1}^{n-1}U_{i}$. Exactness for $(V,U_{n})$ gives $t\in{\mathfrak{F}}(V\wedge U_{n})$ with $\sum_{i=1}^{n-1}\iota_{V,U_{i}}(s_{i})=\iota_{V,V\wedge U_{n}}(t)$ and $s_{n}=-\iota_{U_{n},V\wedge U_{n}}(t)$. By distributivity, $V\wedge U_{n}\cong\bigvee_{i=1}^{n-1}(U_{i}\wedge U_{n})$, hence (by the already proven surjectivity) $t$ lifts to $(t_{i})\in\bigoplus_{i=1}^{n-1}{\mathfrak{F}}(U_{i}\wedge U_{n})$. Subtracting $\alpha((t_{i}))$ reduces to an element in $\ker\beta$ for the (n1)(n-1)-cover of $V$, which lies in $\operatorname{im}\alpha$ by the inductive hypothesis. This proves $\ker\beta=\operatorname{im}\alpha$ and hence the cosheaf axiom for all finite covers. ∎
3.2 DGLA attached to a cover
Let $X$ be a distributive pre-ordered set. Let $W\in X$ and ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ be a cover of $W$. Let $\mathfrak{F}$} be a pre-cosheaf of vector spaces over $X$.
Definition 59.
The Čech chain complex $C_{\bullet}({\mathfrak{U}},W;{\mathfrak{F}})$ is defined as
$C_{n}({\mathfrak{U}},W;{\mathfrak{F}})={\rm Alt}\left[\oplus_{i_{0},\ldots,i_{n}}{\mathfrak{F}}(U_{i_{0}}\wedge\ldots\wedge U_{i_{n}})\right],\quad n\geq 0,$ (104)
where $\rm Alt$} means that we only keep elements that are skew-symmetric under permutations of the indices. The differential $\partial:C_{n}\rightarrow C_{n-1}$ is given by $\partial=\sum_{j=0}^{n}(-1)^{j}\lambda_{j}$, where $\lambda_{j}$ is the canonical map
${\mathfrak{F}}(U_{i_{0}}\wedge\ldots\wedge U_{i_{n}})\to{\mathfrak{F}}(U_{i_{0}}\wedge\ldots\widehat{U_{i_{j}}}\ldots\wedge U_{i_{n}}).$ (105)
Lemma 60.
Let $\mathfrak{F}$} be a coflasque cosheaf of vector spaces over $X$ and $\mathfrak{U}$} be a covering of $W\in X$. The homology of the Čech complex with respect to $\mathfrak{U}$} is 0 for $n>0$ and ${\mathfrak{F}}(W)$ for $n=0$.
Proof.
Let $C^{aug}({\mathfrak{U}},W;{\mathfrak{F}})=\{C_{\bullet}({\mathfrak{U}},W;{\mathfrak{F}})\rightarrow{\mathfrak{F}}(W)\}$ be the augmented Čech complex, where the augmentation $\epsilon:C_{0}({\mathfrak{U}},W;{\mathfrak{F}})=\bigoplus_{i}{\mathfrak{F}}(U_{i})\to{\mathfrak{F}}(W)$ is $\epsilon(({\mathsf{F}}_{i}))=\sum_{i}\iota_{W,U_{i}}({\mathsf{F}}_{i})$.
Proposition 61.
Let $\mathfrak{F}$} be a local Lie system over $X$. Then for any $W\in X$ and any cover $\mathfrak{U}$} of $W$ the 1-shifted augmented Čech complex $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ has a natural structure of a non-negatively graded acyclic DGLA.
Proof.
[…]
Proposition 62.
For a fixed $W\in X$, the map which sends a local Lie system $\mathfrak{F}$} and a cover $\mathfrak{U}$} of $W$ to the DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ is functorial in both $\mathfrak{F}$} and $\mathfrak{U}$}.
Proof.
Functoriality in $\mathfrak{F}$} is clear. To prove functoriality in $\mathfrak{U}$}, note that the refinement map $\phi$ induces a linear map $\hat{\phi}:{\hat{J}}\rightarrow{\hat{I}}$, where $^{\hat{J}}$ is the vector space with basis $\{e_{j}\}_{j\in J}$, hence a linear map $\Lambda\hat{\phi}:\Lambda^{\bullet}{\hat{J}}\rightarrow\Lambda^{\bullet}{\hat{I}}$, hence a linear map $\phi_{*}:{\mathfrak{F}}(W)\otimes\Lambda^{\bullet}{\hat{J}}\rightarrow{\mathfrak{F}}(W)\otimes\Lambda^{\bullet}{\hat{I}}$. One can check that $\phi_{*}$ commutes with the differential and maps the subcomplex $C^{aug}_{\bullet+1}({\mathfrak{V}},W;{\mathfrak{F}})$ to $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$. ∎
We will call $C^{aug}_{\bullet+1}(-,W;-)$ the Čech functor. We will need some variants of the Čech functor. First, we can define a graded local Lie system over a distributive pre-ordered set $(X,\leq)$ in an obvious manner. The construction of the acyclic DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ works in this case as well, except that it may have components in negative degrees.
Second, define a pointed DGLA as a DGLA equipped with a distinguished central cycle of degree 2-2 (which we call the curvature). A morphism of pointed DGLAs is a DGLA morphism which preserves the distinguished central cycle. [^9] The category of pointed DGLAs as defined here is a full subcategory of the category of curved DGLAs as defined in [13]. There, for a curved DGLA with curvature $\mathsf{B}$}, $\mathsf{B}$} is not required to be central and the derivation $\partial$ satisfies $\partial^{2}={\rm ad}_{\mathsf{B}}$.
Definition 63.
A graded local Lie system $\mathfrak{F}$} over $(X,\leq)$ with a terminal object $T\in X$ is called pointed if it is equipped with a distinguished central element ${\mathsf{B}}\in{\mathfrak{F}}(T)$ of degree 2-2. Morphisms in the category of pointed graded local Lie systems are required to preserve the distinguished element.
If $\mathfrak{F}$} is a pointed graded local Lie system, then the DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},T;{\mathfrak{F}})$ is a pointed DGLA. The distinguished central cycle is $\mathsf{B}$} which lies in the augmentation (has Čech degree 1-1). Of course, since the DGLA is acyclic, this central cycle is exact. The construction of a pointed DGLA from a pointed graded local Lie system and a cover of $T$ is functorial in both arguments.
@@ 4 The site of fuzzy semilinear sets
In this section we define and study the site of fuzzy semilinear sets in ${\mathbb{R}}^{n}$. We show that it is equivalent as a site to the pre-ordered set of closed polyhedral subsets of $S^{n-1}$ equipped with a slightly unusual topology. In the following section, we will show that quantum lattice systems give rise to local Lie systems over this site.
@@ 4.1 Semilinear sets and their thickenings
The results of Sections 2 and 3 suggest that we should look for a pre-ordered set of subsets of ${\mathbb{R}}^{n}$ with the following two properties. First, $U\leq V$ whenever there exists $r\geq 0$ such that $U\subseteq V^{r}$. Second, for every two elements $U,V$ there should exist $r\geq 0$ such that $U^{r}$ and $V^{r}$ are transverse. The simplest choice is the set whose elements are finite unions of polyhedra, or equivalently, closed semilinear sets in ${\mathbb{R}}^{n}$.
Definition 64.
A semilinear set in ${\mathbb{R}}^{n}$ is a subset of ${\mathbb{R}}^{n}$ which can be defined by means of a finite number of linear equalities and strict linear inequalities. More precisely, a basic semilinear set in ${\mathbb{R}}^{n}$ is an intersection of a finite number of hyperplanes and open half-spaces, and a semilinear set is a finite union of basic semilinear sets [37].
The set of semilinear sets of ${\mathbb{R}}^{n}$ will be denoted ${\mathcal{S}}_{n}$. Projections ${\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ map ${\mathcal{S}}_{m+n}$ to ${\mathcal{S}}_{m}$ [37]. A map ${\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ is called semilinear iff its graph is a semilinear set of ${\mathbb{R}}^{m+n}$. The composition of two semilinear maps is a semilinear map [37].
@@
Recall that for any set $U\subseteq{\mathbb{R}}^{n}$, we write $U^{r}:=\{x\in{\mathbb{R}}^{n}:\exists y\in U\text{ s.t }d(x,y)\leq r\}$ and call this the $r$-thickening of $U$. It is easy to see that if $U$ is closed, then $U^{r}$ is also closed for any $r\geq 0$.
Lemma 66.
@@
$\Delta_{r}=\{(x,y)\in{\mathbb{R}}^{2n}|d(x,y)\leq r\}.$ (106)
By the previous lemma, $\Delta_{r}$ is semilinear. On the other hand, $U^{r}$ is the projection to the first ${\mathbb{R}}^{n}$ of $\Delta_{r}\cap({\mathbb{R}}^{n}\times U)\subseteq{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}$. Since intersection and projection preserve the set of semilinear sets, the lemma is proved. ∎
Lemma 67.
@@
Suppose $x,y\in U^{r}$, and suppose $x^{\prime},y^{\prime}\in U$ satisfy $d(x,x^{\prime})\leq r$ and $d(y,y^{\prime})\leq r$. Then for any $t\in[0,1]$ we have $d(tx+(1-t)y,tx^{\prime}+(1-t)y^{\prime})\leq td(x,x^{\prime})+(1-t)d(y,y^{\prime})\leq r.$ Since $U$ is convex, $tx^{\prime}+(1-t)y^{\prime}\in U$, hence $tx+(1-t)y\in U^{r}$. ∎
A polyhedron in ${\mathbb{R}}^{n}$ is an intersection of a finite number of closed half-spaces. A polyhedron is closed, but not necessarily compact. A closed semilinear set is the same as a finite union of polyhedra. Conversely, according to Theorem 19.6 of [33], a polyhedron can be described as a closed convex semilinear set. Combining this with the above lemmas, we get
Corollary 68.